OpenStudy (anonymous):

Is the set of all 2x2 matrices of the form (row1:[a b] row2:[a+b 0]) a subspace of the set of all 2x2 matrices?

OpenStudy (kc_kennylau):

Welcome to OpenStudy :D You may want to read the code of conduct ( http://openstudy.com/code-of-conduct ). Don't forget to click on the button called "Best Answer" to give a medal to the person who helped you :)

OpenStudy (anonymous):

like this? \[M= \left[ \begin{array}{c} a & b \\ a+b & 0 \end{array} \right] \] (( reference /use \[\text{\[M= \left[ \begin{array}{c} a & b \\ a+b & 0 \end{array} \right] \] } \] ))

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hard to write like that :P

OpenStudy (zarkon):

no

OpenStudy (kc_kennylau):

Rephrased code: \[\mbox{\[M=}\color{blue}{\mbox{\left[}}\color{green}{\mbox{\begin{array}{c}}}\mbox{ a & b \\ a+b & 0 }\color{green}{\mbox{\end{array}}} \color{blue}{\mbox{\right]}}\mbox{\] }\]

OpenStudy (kc_kennylau):

i mean coloured*

OpenStudy (anonymous):

:'( Well darn.

OpenStudy (zarkon):

addition is coordinate wise

OpenStudy (zarkon):

\[\left[\begin{array}{cc} \alpha & \beta \\ \gamma & \delta \\ \end{array}\right]+\left[\begin{array}{cc} \epsilon & \zeta \\ \eta & \theta \\ \end{array}\right]=\left[\begin{array}{cc} \alpha+\epsilon & \beta+\zeta \\ \gamma+\eta & \delta+\theta \\ \end{array}\right]\]

OpenStudy (anonymous):

OH! thanks. Brain completely jumbled that up.

OpenStudy (anonymous):

\[M_1 + M_2 = \left[ \begin{array}{c} a_1 & b_1 \\ a_1+b_1 & 0 \end{array} \right] + \left[ \begin{array}{c} a_2 & b_2 \\ a_2+b_2 & 0 \end{array} \right] \] \[ = \left[ \begin{array}{c} (a_1+a_2) & (b_1+b_2) \\ (a_1+b_1)+(a_2+b_2) & 0\end{array} \right] \] So that does work. And so will multiplication. also \[M= a\left[ \begin{array}{c} 1 & 0 \\ 1 & 0 \end{array} \right] +b\left[ \begin{array}{c} 0 & 1 \\ 1 & 0 \end{array} \right]\] so it's a subspace?

OpenStudy (anonymous):

~vector multiplication \[rM= r\left[ \begin{array}{c} a & b \\ a+b & 0 \end{array} \right]= \left[ \begin{array}{c} ra & rb \\ r(a+b) & 0 \end{array} \right] \]

OpenStudy (anonymous):

those all seem like trivial things to do, however. I was just basing this off another proof I found, but it doesn't seem particularly proof-y....

OpenStudy (zarkon):

\[=\left[\begin{array}{cc} ra & rb \\ ra+rb & 0\\ \end{array}\right]\] now it is of the form as the original matrix should probably include this little step

OpenStudy (zarkon):

also, one should show that you can get the zero vector (zero matrix in this case)

OpenStudy (anonymous):

Does the ability to separate it into \[M= a\left[ \begin{array}{c} 1 & 0 \\ 1 & 0 \end{array} \right]+b\left[ \begin{array}{c} 0 & 1 \\ 1 & 0 \end{array} \right]\] automatically show that it's a subspace? Or do you have to show those other operations as well? It almost seems more like you're defining those properties onto the matrices more than you're proving that it works - or is it just that this is a simple example? Lastly, \[ 0M= 0\left[ \begin{array}{c} a & b \\ a+b & 0 \end{array} \right]=\left[ \begin{array}{c} 0(a) & 0(b) \\ 0(a+b) & 0 \end{array} \right]=\left[ \begin{array}{c} 0 & 0 \\ 0(a)+0(b) & 0 \end{array} \right]=\left[ \begin{array}{c} 0 & 0 \\ 0 & 0 \end{array} \right]\] ?

OpenStudy (anonymous):

That last bit is wrong methinks. That's just the multiplication. Would you have to define a=a+b=b =0 ?

OpenStudy (zarkon):

for your first question...No you are just showing that your matrix can be written as a linear combination of those two vectors. That does not show it is a subspace

OpenStudy (zarkon):

right...just let a=b=0 and you get the zero vector

OpenStudy (anonymous):

Ok, thanks! So, for instance, if it were \[M= \left[ \begin{array}{c} a+1 & b \\ a+b & 0 \end{array} \right]\] since the zero matrix isn't in M, then that would not a subspace of the set of all 2x2 matrices?

OpenStudy (anonymous):

Thanks for all the help, @Zarkon ^_^

OpenStudy (phi):

** Is the set W of all 2x2 matrices of the form (row1:[a b] row2:[a+b 0]) a subspace of the set of all 2x2 matrices? ** I would say yes. W is a subspace iff The zero vector, 0, is in W. If u and v are elements of W, then the sum u + v is an element of W; If u is an element of W and c is a scalar from K, then the product cu is an element of W; you have the zero matrix, (set a = b= 0) and, as posted above, any linear combination of u, and v (both in W) result in a matrix contained in W.

OpenStudy (zarkon):

@AllTehMaffs right...that set \(M\) you just made does not contain the zero vector so it can't be a subspace